Assume that in 2015 the United States will need 3.0*10^12watts of electric power produced by thousands of 1000 MW power plants. These will almost certainly be powered by either coal or nuclear reactors. If all the energy is provided by U-235 burning reactors, how many tonnes of U-235 will be needed annually? NOTE: assume that fission of U-235 releases a "rounded" 8.E13 J/kg (Assume a realistic efficiency of 40 percent for the power plant)
Answer #1

power consumption = power produced

3.0*10^12watts = m kg/sec * 8.0E13 J/kg * 0.4 = mass * energy per kg *efficiency

m = 3* 10^12 / (8 * 10^13 *.4 )

m = 3/(8*4) kg/sec

masss in a year = 3*60*60*24*365.25/(8*4)

= 2958525 Kg /year